# Language-centered Human Activity Recognition

### v2

@@ 5. Discussion
Lessons learned: We summarized the following lessons to benefit future studies: (1) Introducing semantic interpretation can mitigate heterogeneity in cross-dataset HAR and offers the potential for recognizing new activities, as indicated in Table 2 and Table 3. (2) The success of LLMs in cross-dataset HAR is partly due to their ability to summarize patterns and represent them using language. This leads to lower KL divergence between cross-dataset data distributions compared to raw data. (3) Improving the quality of LLM responses through iterative re-generation, guided by KL divergence, enhances HAR performance. The evaluation shows that generating high-quality responses has a direct positive impact on HAR accuracy.
Comparison to other ways of utilizing LLMs: In addition to our approach, there are two other potential methods for utilizing LLMs in HAR: training a foundation model specifically for HAR or fine-tuning an existing open-source LLM, such as Llama. A key advantage of our approach lies in its low cost and flexibility. On one hand, our method does not require large amounts of data or computing resources for model training or fine-tuning. On the other hand, different LLMs can be flexibly integrated into our framework. As demonstrated in our evaluation, incorporating a high-quality LLM could further enhance the performance of our method.
Other ways of improving LLMs responses: There are ongoing efforts in the AI community aimed at improving the trustworthiness of LLMs through methods such as retrieval-augmented generation and memory-augmented models. Our approach offers a general framework for post-hoc factuality verification, where advancements in these AI methods can be seamlessly integrated into our framework.
Limitations: The process of obtaining semantic interpretations using LLMs is time-consuming, primarily due to the slow inference speed of current LLMs. We anticipate that future advancements in accelerating LLMs will enhance the efficiency of our approach. Although we used four public datasets in the evaluation, they remain relatively small in scale with a limited number of activities. A larger dataset could provide a more comprehensive evaluation and further validate our method’s effectiveness.
@@ 6. Related work
Human activity recognition: Human Activity Recognition (HAR) leverages data from various sensors such as accelerometers and gyroscopes to detect and classify human activities (Yao et al., 2017). To mitigate the issue of data heterogeneity, existing work can be classified into three categories, including self-supervised learning methods, domain adaptation methods, and data augmentation methods. Self-supervised methods (Xu et al., 2021; Saeed et al., 2019) aim to maximize the use of unlabeled data to extract valuable features. This approach allows models to require only a minimal amount of labeled data to efficiently adapt to downstream tasks. Despite their capabilities, these models still depend on a certain amount of labeled data to train classifiers, which may underperform when applied to target datasets that lack any labeled data. The core idea of domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) involves taking a classifier that has been pre-trained on a source domain and using it with an unlabeled dataset from a target domain, then adjusting the source model’s weights to enhance its performance in the target domain. Data augmentation (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019; Wang et al., 2022; Tang et al., 2021) in HAR aims to increase the diversity and quality of training data, thereby enhancing the model’s generalization ability and accuracy. However, they still face data heterogeneity and heterogeneity can amplify the impact of classification errors on downstream applications. Additionally, cross-dataset scenarios often introduce new activities that existing methods struggle to handle.
@@
Some work (Cheng et al., 2013) attempts to address the challenge of recognizing new activities by using predefined semantics, such as actions of specific body parts. These approaches train two mappings: one from sensor readings to predefined semantics, and another from predefined semantics to activity labels. However, the quality of these predefined semantics heavily impacts the model’s performance. Additionally, this work are not well-suited for cross-dataset new activities, as the new activities may differ significantly from the original dataset.
@@
LLM for human activity recognition: Recently, there has been a surge of interest in leveraging the capabilities of large language models (LLMs) to enhance mobile computing applications (Leng et al., 2024). UniHAR (Xu et al., 2023) explore the potential of leveraging large language models (LLMs) to comprehend and interact with the physical world. Kim et al. (Kim et al., 2024) harness the power of large language models (LLMs) to improve health prediction using data collected from wearable sensors. IMUGPT2.0 (Leng et al., 2024) utilize LLMs to help data generation since IMU-labeled data is challenging to collect. MotionGPT (Jiang et al., 2023) fuse language data with large-scale motion models, and motion-language pre-training that can enhance the performance of motion-related tasks becomes feasible.

### v3

@@ 5. Discussion
Lessons learned: We summarize the following lessons to benefit future studies: (1) Introducing semantic interpretation can mitigate heterogeneity in cross-dataset HAR and offers the potential to recognize new activities, as indicated in Table 2 and Fig. 10. (2) The success of LLMs in cross-dataset HAR is partly due to their ability to summarize patterns and represent them using language. This leads to a lower KL divergence between cross-dataset data distributions compared to raw data. (3) Improving the quality of LLM responses through iterative re-generation, guided by the KL divergence, enhances HAR performance. The evaluation shows that generating high-quality responses has a direct positive impact on HAR accuracy.
Comparison to other ways of utilizing LLMs: In addition to our method, there are two other potential methods for utilizing LLMs in HAR: training a foundation model specifically for HAR or fine-tuning an existing open-source LLM, such as Llama. A key advantage of our method lies in its low cost and flexibility. On the one hand, our method does not require large amounts of data or computing resources for model training or fine-tuning. On the other hand, different LLMs can be flexibly integrated into our framework. As demonstrated in our evaluation, incorporating a high-quality LLM could further enhance the performance of our method.
Other ways of improving LLMs responses: There are ongoing efforts in the AI community aimed at improving the trustworthiness of LLMs through methods such as retrieval-augmented generation and memory-augmented models. Our method offers a general framework for post-hoc factuality verification, where advancements in these AI methods can be seamlessly integrated into our framework.
Limitations: The process of obtaining semantic interpretations using LLMs is time-consuming, primarily due to the slow inference speed of current LLMs. We anticipate that future advancements in accelerating LLMs will enhance the efficiency of our method. Although we used four public datasets in the evaluation, they remain relatively small in scale with a limited number of activities. A larger dataset could provide a more comprehensive evaluation and further validate our method’s effectiveness.
@@ 6. Related work
Human activity recognition: Human Activity Recognition (HAR) leverages data from various sensors such as accelerometers and gyroscopes to detect and classify human activities (Yao et al., 2017). To mitigate the issue of data heterogeneity, existing works can be classified into three categories, including self-supervised learning methods, domain adaptation methods, and data augmentation methods. Self-supervised methods (Xu et al., 2021; Saeed et al., 2019) aim to maximize the use of unlabeled data to extract valuable features. This method allows models to require only a minimal amount of labeled data to efficiently adapt to downstream tasks. Despite their capabilities, these models still depend on a certain amount of labeled data to train classifiers, which may struggle when applied to target datasets that lack any labeled data. The core idea of domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) involves taking a classifier that has been pre-trained on a source domain and using it with an unlabeled dataset from a target domain, then adjusting the source model’s weights to enhance its performance in the target domain. Data augmentation (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019; Wang et al., 2022; Tang et al., 2021) in HAR aims to increase the diversity and quality of training data, thereby enhancing the model’s generalization ability and accuracy. However, these methods still face data heterogeneity and heterogeneity can amplify the impact of classification errors on downstream applications. Additionally, cross-dataset scenarios often introduce new activities that existing methods struggle to handle.
@@
Some works (Cheng et al., 2013) attempts to address the challenge of recognizing new activities by using pre-defined semantics, such as actions of specific body parts. These methods train two mappings: one from sensor readings to predefined semantics, and another from predefined semantics to activity labels. However, the quality of these predefined semantics significantly influences the model performance. Additionally, this work is not well-suited for cross-dataset new activities, as the new activities may differ significantly from the original dataset.
@@
LLM for human activity recognition: Recently, there has been a surge of interest in leveraging the capabilities of large language models (LLMs) to enhance mobile computing applications (Leng et al., 2024). UniHAR (Xu et al., 2023) explores the potential of leveraging large language models (LLMs) to comprehend and interact with the physical world. Kim et al. (Kim et al., 2024) harness the power of large language models (LLMs) to improve health prediction using data collected from wearable sensors. IMUGPT2.0 (Leng et al., 2024) utilizes LLMs to help data generation since IMU-labeled data is challenging to collect. MotionGPT (Jiang et al., 2023) fuses language data with large-scale motion models, and motion-language pre-training that can enhance the performance of motion-related tasks becomes feasible.